Determine the equation of the line that has a distance 2 from the origin and is
parallel to the line that passes thru P1(2,7) and P2(-3,-2).
1
Expert's answer
2020-09-29T18:01:26-0400

Distance of line from origin d=2

Since the required line is parallel to the line formed by joining "P_1(2,7) and P_2(-3,-2)" ,

There slope must be equal

Let m be the slope of required line

"m=\frac{-2-7}{-3-2}"

"=\frac{-9}{-5}"

so m="\frac{9}{5}"

Let the equation of required line be

"y=mx+c"

"y=\frac{9}{5}x+c"

Rearranging the equation as,

"9x-5y+5c=0"


Distance from origin is given by,

d="\frac{|9\times 0+0\times -5+5c|}{\sqrt{9^2+(-5)^2}}"


2="\frac{5c}{\sqrt{106}}"


5c="2\sqrt{106}"


c=0.4"\sqrt{106}"

So the equation of required line is


9x-5y+2"\sqrt{106}" =0